### v1

# BOUQuET [Uncaptioned image]: dataset, Benchmark and Open initiative for Universal Quality Evaluation in Translation

Abstract
This paper presents BOUQuET, a multicentric and multi-register/domain dataset and benchmark, and its broader collaborative extension initiative. This dataset is handcrafted in non-English languages first, each of these source languages being represented among the 23 languages commonly used by half of the world’s population and therefore having the potential to serve as pivot languages that will enable more accurate translations. The dataset is specially designed to avoid contamination and be multicentric, so as to enforce representation of multilingual language features. In addition, the dataset goes beyond the sentence level, as it is organized in paragraphs of various lengths. Compared with related machine translation (MT) datasets, we show that BOUQuET has a broader representation of domains while simplifying the translation task for non-experts. Therefore, BOUQuET is specially suitable for the open initiative and call for translation participation that we are launching to extend it to a multi-way parallel corpus to any written language.
1 Introduction
Although multilingual large-language model evaluation benchmarks are only starting (Dac Lai et al., 2023), there is a rich research history in multilingual evaluation datasets for natural language processing; e.g., Sun and Duh (2020); Malmasi et al. (2022); Yu et al. (2022), with Machine Translation (MT) being the task with the highest investment in multilinguality (Kocmi et al., 2024). This is evident from the nearly 15-year history of the Conference on Machine Translation (formerly a workshop, WMT), which has established an international evaluation campaign (Kocmi et al., 2024). The campaign has compiled a comprehensive collection of parallel data evaluations covering a broad range of language pairs, domains, and tasks. Although some datasets offer multi-way parallelism across languages (e.g., the TED parallel corpus), [^1] https://hal.science/hal-04702210v1 the largest multi-way parallel corpus was only recently introduced with FLORES-101 (Goyal et al., 2022), later expanded to FLORES-200 (NLLB-Team et al., 2022) and to 2M-FLORES (Costa-jussà et al., 2024).
However, even if great progress has been made, advances in multilingual applications for diverse languages are hindered by limited evaluations. Existing datasets and benchmarks fall short due to having an English-centric focus, a narrow register, compromised quality from automated construction and mining, limited language coverage, or a static nature, in addition to being prone to contamination.
Similarly, in parallel with the previous progress, there have been several initiatives that called for data annotation in a collaborative and open way, such as the translation data collection initiative reported in Singh et al. (2024).
This precisely sets the stage for the introduction of a multilingual evaluation dataset and a benchmark, BOUQuET, which additionally combines community efforts. The purpose of this paper is two-fold. First, the paper details how we develop the Source-BOUQuET dataset, which is the necessary stepping stone towards an open initiative. Second, it presents how we design the open initiative itself, which aims to build the Full-BOUQuET dataset; i.e., Source-BOUQuET translated into any written language.
1.1 Definitions
Before describing the Source-BOUQuET dataset’s characteristics and building methodology, we define our use of some frequently encountered terms that may cover a variety of meanings.
Domain.
By the term domain, we mean different spaces in which language is produced in speech, sign, or writing (e.g., books, social media, news, Wikipedia, organization websites, official documents, direct messaging, texting). In this paper, we focus solely on the written modality.
Register.
We understand the term register as a functional variety of language that includes socio-semiotic properties, as expressed in Halliday and Matthiessen (2004), or more simply as a “contextual style”, as presented in Labov , pp.79–99). In that regard, a register is a specific variety of language used to best fit a specific communicative purpose in a specific situation.
2 Dataset: Source-BOUQuET
In this section, we describe the creation criteria that have follow to design Source-BOUQuET, as well as the languages it includes.
2.1 Main characteristics
As described in greater detail next, the Source-BOUQuET dataset is mainly characterized by its non-English-centric focus, its diverse registers and domain (which is complementary to FLORES-200), its manual and original composition, and its built-in dynamic extensibility.
Non-English-centric focus.
Source-BOUQuET is handcrafted by proficient speakers of French, German, Hindi, Indonesian, Mandarin Chinese, Russian, and Spanish. Each of those languages provides the same number of sentences to the final dataset. The languages for Source-BOUQuET (see Table 2 in Section 2.3) are all part of the top 20 languages in the world in terms user population, as listed in Eberhard et al. (2024). In addition, they are also used by a large number of both native and non-native speakers, which makes them good candidates for what we refer to as pivot languages; i.e., higher-resource languages that can facilitate—as source languages—the translation of datasets into lower-resource languages. English is often used as such a pivot language, since numerous people have a high degree of proficiency in English as a second language. English is not the only language in this situation, however, and is not always the best pivot language option. For example, it is much easier to find Guarani-Spanish bilingual speakers than it is to find Guarani-English bilingual speakers. What is more, cultural proximity may also make translation slightly easier.
Diverse registers and domains.
Registers derive from communicative purposes and, as such, are related to domains. However, the relationship between registers and domains is not one to one. For example, if we take a domain such as TV news, we can identify at least 3 registers:
  • The register used by the news anchor, which is represented by fully scripted language that is read from a teleprompter with a very specific and unnatural form of diction (hypercorrect enunciation, unnatural intonation, homogeneous pace).
  • The register produced by communication specialists (i.e., people who have been trained to be spokespersons or surrogates). The points they make have been scriptedand rehearsed to the point of being known by heart. Communication specialists also learn rehearsed patterns of language to make sure they do not get surprised by a question and can redirect the conversation towards the points they need to make instead of answering unexpected questions. All the language they use has been written and rehearsed. It sounds spontaneous but it is not structured like informal language.
  • The register represented in person-in-the-street segments, which is more informal and spontaneous (possibly colloquial).
The example above is taken from a domain where both speech and writing are used but the situation is not significantly different in the written modality only. Language users all commonly shift between registers, which is typically referred to as style-shifting. Style-shifting (i.e., register-shifting) occurs within domains; so the domain itself is not a fool-proof way of getting a specific register. Although the norms of the domain can impose the degree of formality and of lexical specialization, it is often the register (which derives from the communicative purpose), not the domain, that determines many aspects of linguistic structure (lexical density, pronoun use, syntax, etc.).
The registers we cover is Source-BOUQuET are given in Figure 1.
Manual construction and original composition (not crawled) with accurate revisions
To develop Source-BOUQuET, we set a variety of linguistic criteria that need to be covered. Guidelines are then shared with linguists who manually craft sentences covering examples of these linguistic criteria and compose paragraphs of various lengths.
The main strategies for open collaboration are to design contribution guidelines and build the annotation tool that enables to freely collect translations in any language. For this, BOUQuET is initially composed on 8 major languages plus the English translation and it is shared in a repository that allows language communities to easily add a new language even if they do not speak English, e.g. a native speaker of Quechua can translate from Spanish. This repository contains detailed guidelines on how to do it. BOUQuET’s innovative approach ensures universal language accessibility. This open collaborative initiative will enrich BOUQuET with the following characteristics.
Language coverage extensibility
Using both private and community-driven initiatives, we could potentially support any written language, as long as there is individual interest in contributing to multilingual advancements.
Dynamic in nature
Since BOUQuET includes the community, it can continuously evolve by constantly engaging it.
2.2 Creation criteria
For the design of the creation guidelines, which are reported in detail in Appendix A, we prepared a list of linguistic coverage requirements as well as some statistical information.
Linguistic coverage requirements.
In order for BOUQuET to be representative of various linguistic phenomena, linguistic coverage requirements are defined (as listed in Table 1), which are to be included in sentences that form paragraphs. This continuity will include variation in the number of sentences per paragraph and in the lengths of sentences themselves.
Phenomena
Paragraph-like continuity
Variation in sentence lengths
Dominant (unmarked) and non-dominant (marked) word orders
Different emphasis or topicalization
Different sentence structures (affirmation, interrogation, negation, subordination, coordination)
Different verb moods, tenses, and aspects
Different morphosyntactic options
Different grammatical persons (1st, 2nd, 3rd, singular, plural)
Different grammatical genders
Different grammatical number agreement
Different grammatical case or forms of inflection
Most frequent words used in various registers
Presence of named entities, numbers, slang, and emojis
Table 1: Source-BOUQuET Linguistic Requirements
Variety of domains.
Source-BOUQuET is intended to cover 8 domains: narration (as in fiction writing), dialog, social media posts, social media comments, how-to, miscellaneous website content (excluding social media or news), opinion pieces, miscellaneous other (such as written speeches or signage). The choice of these domains optimizes for variety and popular usefulness.
Variety of registers.
Source-BOUQuET is built with register variety in mind, differently from FLORES-200, which covers a few different domains but remains largely within similar registers. We characterize the registers through 3 main aspects (connectedness, preparedness, and social differential), and a variety of options within these aspects (see Figure 1 for detail). Connectedness attempts to describe the type of interaction typically available in a given domain. Preparedness aims to gauge how much time is typically used to produce or edit language content. Social differential tries to describe various social situations in which language content is typically produced in each of the domains. Options are not mutually exclusive at the level of the domain but become differentiated at the level of a single sentence.
By including new registers and domains, the new dataset is likely to be more generalizable to different contexts and applications.
Figure 1: Covered register characteristics.
Statistical guidance.
In order to appropriately cover linguistic requirements and adequately represent the registers and domains, we performed a statistical analysis to understand each of the linguistic characteristics. In particular, our analysis covers most domains that we are including in Source-BOUQuET by using diverse public datasets: narration (Books3, gutenberg (Gerlach and Font-Clos, 2018)); Social media posts (Reddit (Baumgartner et al., 2020)); Social media comments (Wikipedia comments [^2] https://www.kaggle.com/competitions/jigsaw-multilingual-toxic-comment-classification); Conversations / Dialogues (dialogsum (Chen et al., 2021), Open Orca (Lian et al., 2023)); Tutorials/how-to articles (how-to Wikipedia-lingua  [^3] https://huggingface.co/datasets/GEM/wiki_lingua); Website content (c4 (Raffel et al., 2020)); News / Reflection pieces (cnndailymail (Nallapati et al., 2016), xsum (Narayan et al., 2018)) and Miscellaneous (Wikipedia). Note that we collect information from public data that do not always accurately match our categories but constitute a proxy. For each of these domains, we have analyzed dimensionality: characters per token; tokens per sentence and sentences per paragraph.
Figure 2: (Left) Tokens per sentence and (middle) sentences per paragraph (Right) CEFR per domain.
Regarding tokens per sentence (Figure 2 left), we can see correlations between different domains, and clear differences in length, especially in dialogs which tend to be much shorter.
Regarding sentences per paragraph (Figure 2 middle), we can find correlation between different datasets representing the same domain, where fiction writing paragraphs tend to be much longer (averaging 5 but reaching up to 20 ), dialogs and news articles are much shorter (barely reaching 3-4 sentences in a paragraph), and the rest of the categories are somewhere in between (normally staying between 1-5 but reaching up to 10 in some cases).
Furthermore, we have checked complexity: distribution of CEFR level [^4] https://rm.coe.int/common-european-framework-of-reference-for-languages-learning-teaching/16809ea0d4 as a proxy. Including here the % of C2 scores at sentence level for each dataset. This was labeled by a sonar-based model (Duquenne et al., 2023) trained on CEFR-SP data CEFR Text Classification which greatly outperformed Llama-3.1 (Touvron et al., 2023). See Figure 2 right.
Wikipedia seems to be the only dataset with a more considerable share of C2 sentences, with some others like dialogues having no samples scored as such.
Annotations.
Each entry of Source-BOUQuET includes the source text (in one of the 8 pivot languages from Table 2) and its translation into English, domain information, and contextual information for better translation accuracy.
2.3 Languages
BOUQuET aims to be as multicentric as possible, in contrast with most existing datasets like FLORES-200 that are English-centric. The motivation is mainly to be representative of linguistic phenomena. To this effect, it is created in 7 non-English languages. Each language contributes with a similar number of sentences along with their English equivalents given by the sentence creators themselves. See languages in Table 2.
ISO 639-3 ISO 15924 Language Family Subgroup1
cmn Hans Mandarin Chinese Sino-Tibetan Sinitic
deu Latn German Indo-European Germanic
fra Latn French Indo-European Italic
hin Deva Hindi Indo-European Indo-Aryan
ind Latn Indonesian Austronesian Malayo-Polynesian
rus Cyrl Russian Indo-European Balto-Slavic
spa Latn Spanish Indo-European Italic
Table 2: Source-BOUQuET Languages
3 Benchmark
Statistics
In total, BOUQuET currently contains 1,750 sentences. These sentences are split by doing a stratified selection among source languages and domains into development and test sets with an equivalent number of sentences. A portion of the development set (522) is intended to initially be kept hidden. Results in the following sections are presented with the test split of 889 sentences.
Table 3 provides a comparison of several relevant statistics from BOUQuET and the closest related datasets including FLORES-200, NTREX-128 (Federmann et al., 2022) and NLLB-MD (NLLB-Team et al., 2022).
Dataset Sent. Parag. Reg. Dom. Lang. Dyn.
FLORES-200 3,003 () $\times$× Wikipedia, News, Travel guides 200 $\times$×
NTREX-128 1,997 $\times$× News 128 $\times$×
NLLB-MD 9,000 $\times$× $\times$× Chat, News, Health 6 $\times$×
BOUQuET 1,750 Fiction, Conversation, Social media posts/comments 8+
Tutorials, Website, Reflection pieces, Miscellaneous
Table 3: Main statistics from MT evaluation datasets including BOUQuET: number of sentences, paragraph-information, registers, domains, languages, dynamism.
Domain representation
The performance of the model in a new or unseen dataset depends on the similarity between the dataset that was used to fit the model and the new dataset. In order to more broadly evaluate an MT model, we are interested in covering a different set of domains. We compare the domain coverage of BOUQuET with that of FLORES-200, NTREX-128 (Federmann et al., 2022) and NLLB-MD (Stolte et al., 2024) . To do this comparison, we represent data (multi-domains public datasets from Figure 2 in grey, each alternative dataset FLORES-200, NTREX-128 and NLLB-MD and BOUQuET) with SONAR (Duquenne et al., 2023). From SONAR vectors, we do a PCA-dimensionality reduction, see Figure 3. All domains from Figure 2 are represented in grey. Then, Figure 3 from left to right compares BOUQuET against FLORES-200, NTREX-128 and NLLB-MD, respectively. We qualitatively observe that BOUQuET covers a wider range of domains.
Additionally, to quantify this coverage, we measure the overlap between each dataset with each of the domains using two metrics, the Wasserstein distance (implemented with POT library [^5] https://pythonot.github.io/) and the Silhouette coefficient (implemented with sklearn [^6] https://scikit-learn.org/stable/). The Wasserstein Distance (WD), also known as the Earth Mover’s Distance (EMD), is a metric that measures the "effort" required to transform one probability distribution into another. Lower results indicate a higher similarity between clusters. The lowest consistent results for all domains are obtained with the data set BOUQuET. Alternatively, the Silhouette Score ($S$)  [^7] https://en.wikipedia.org/wiki/Silhouette_(clustering) measures how well each data point fits within its assigned cluster compared to other clusters, it ranges between -1 and 1 and the higher the coefficient, the more coherent the clusters are. Again, BOUQuET shows a lower score, consistent across domains, meaning closer representation across domains.
Figure 3: Domain representation and overlap with datasets (FLORES-200, NTREX-128, NLLB-MD)
Narration Website Opinion Dialog Social media Social media How-to Misc.
comments posts
WD SC WD SC WD SC WD SC WD SC WD SC WD SC WD SC
FLORES-200 0.32 0.05 0.32 0.03 0.32 0.02 0.31 0.11 0.32 0.09 0.31 0.08 0.31 0.06 0.33 0.03
NTREX-128 0.32 0.04 0.31 0.03 0.32 0.01 0.31 0.09 0.31 0.07 0.31 0.6 0.31 0.06 0.33 0.03
NLLB-MD 0.31 0.04 0.31 0.02 0.31 0.00 0.30 0.08 0.31 0.06 0.30 0.05 0.30 0.05 0.32 0.02
Bouquet 0.29 -0.02 0.29 -0.02 0.30 -0.03 0.28 0.02 0.29 0.01 0.28 0.00 0.28 0.00 0.30 -0.01
Table 4: Wasserstein Distance (WD) and Silhouette Score ($S$) for each domain and dataset. Lower WD and $S$ indicate better representation of the domain. Best results in bold.
MT Benchmarking
To help the reader understand why the dataset is useful, we present preliminary results to demonstrate its use for its intended purpose: MT benchmarking. We evaluated on translation systems (GPT-4o (Wu et al., 2024), Llama-3.1 (Llama-3.3-8B)(Touvron et al., 2023), and NLLB (NLLB-3.3B) (NLLB-Team et al., 2022)) using two metrics: Comet (CometKiwi-da-xl, range 0-1 and $\uparrow$ better) (Chimoto and Bassett, 2022) and MetricX (MetricX-24, range 0-25 and $\downarrow$ better) (Juraska et al., 2024), following WMT 2024 (Kocmi et al., 2024). See results in Table 5. We observe that BOUQuET scores consistently higher than NLLB-MD on average and for all models, suggesting that BOUQuET is easier to translate. This is an advantage for the open initiative, since the complexity of current MT test sets makes it harder to ask the community to participate in translations as it requires a high-level of expertise. Rankings across models nor languages is not preserved, which hints that BOUQuET may be posing different challenges to the models. We need to further investigate which linguistic challenges BOUQuET is adding. On this ranking, if we look at average BOUQuET, LlaMa is the worst system according to Comet but the best system according to Metricx. This does not happen in NLLB-MD, where the ranking is consistent across metrics. If we look per language, ranking across metrics is consistent for 4 out of 7 languages for NLLB-MD and for 1 out of 7 languages for BOUQuET. This raises a concern and may suggest that MT metrics themselves do not generalize well to the broader domains/registers in BOUQuET. However, this requires further investigation and we want to specifically explore different alternatives to use BOUQuET such as exploring the performance by domain and register. Note that we present only preliminary experiments, see Section 5 for more details on ongoing experiments.
M cmn deu fra hin ind rus spa AVG
COM($\uparrow$) MetX($\downarrow$) COM MetX COM MetX COM MetX COM MetX COM MetX COM MetX COM MetX
NLLB-MD
GPT4o 0.76 14.34 0.77 6.80 0.76 13.86 0.68 6.63 0.79 16.16 0.78 3.96 0.80 13.99 0.76 10.81
Llama 0.71 14.38 0.71 6.96 0.72 14.00 0.57 7.07 0.75 16.21 0.71 4.00 0.75 14.11 0.70 10.96
NLLB 0.61 14.80 0.70 7.01 0.70 13.98 0.64 6.72 0.73 16.20 0.72 4.43 0.76 14.06 0.69 11.03
BOUQuET
GPT4o 0.79 1.99 0.80 2.79 0.78 5.57 0.76 3.37 0.84 4.07 0.81 3.60 0.83 4.74 0.80 3.73
Llama 0.75 2.06 0.76 2.07 0.74 4.68 0.72 3.19 0.83 3.54 0.75 3.26 0.80 4.07 0.76 3.27
NLLB 0.67 2.19 0.79 2.44 0.77 5.30 0.74 3.49 0.82 3.94 0.78 3.40 0.67 4.42 0.77 3.60
Table 5: Results on Quality Estimation from English to 7 Source-BOUQuET languages on BOUQuET evaluation dataset .
4 Open initiative
Source-BOUQuET is intended to be translated into any written language. This ambition can only be achieved with the support of the community. For this, we have organized an open collaborative effort which involves language communities that are interested in contributing to this effort.
The purpose of this open initiative is to collect translations from Source-BOUQuET. To collect these annotations, we have chosen the ARGILLA tool [^8] https://argilla.io/ which is a collaboration platform for building datasets, designed for use by AI engineers and domain experts.
Together with setting the Source-BOUQuET in this tool, we have designed annotation guidelines which very much resemble those from FLORES-200 (NLLB-Team et al., 2022).
This open initiative is available in HuggingFace https://huggingface.co/spaces/facebook/bouquet and it is starting as part of the Meta Language Technology Partner Program [^9] https://docs.google.com/forms/d/e/1FAIpQLSdzcRdtkQCuTrXw727DgJgWbOPKDj5v0bArgGfQUTT6sEopFw/viewform publicly launched in AI Summit [^10] https://www.elysee.fr/en/sommet-pour-l-action-sur-l-ia.
5 Conclusions and Next Steps
In this paper, we have presented the Source-BOUQuET dataset and the attached open initiative. This dataset is available in 8 languages. We have shown that consistent BOUQuET gains in domain diversity by in two different metrics while keeping complexity lower than its competitors. The latter is specially relevant to simplify the translation for non-experts that may join the open initiative. This paper and Source-BOUQuET is only a fraction of the ambitious project that we are starting by launching this open call for community efforts. Please join us in making Universal Quality Evaluation in Translation available in any language!
Limitations
BOUQuET is work in progress; the BOUQuET dataset is still limited in number of languages and translations. The benchmarking is also limited in several axes (dataset comparison, models and metrics evaluated). However, we prioritized starting the BOUQuET open initiative in coordination with the Meta Language Technology Partner Program launch so that we could start raising the collaboration interest. As OmnilingualMT team, we continue actively investing on it, see next steps that we are planning to release soon.
Coming Soon
BOUQuET is actively evolving and we are currently working on making the following progress:
  • Adding 250 sentences in Source-BOUQuET in Egyptian Arabic.
  • Complete Source-BOUQuET in all 7 languages from Table 2 plus Egyptian Arabic.
  • The creation guidelines plus translation guidelines in all 8 pivot languages.
  • Designing quality control for each of the contributions and adding new languages to the incremental releases of BOUQuET.
  • Extending the Benchmarking by further showing the capabilities of BOUQuET, e.g. increasing the evaluation of linguistic signals over its alternatives (FLORES-200, NTREX-128, and NLLB-MD). For this, we propose to explore the performance in different domains and see if we get difference in the performance ranking of translation systems.
Acknowledgments
We would like to extend our sincere gratitude to Luis Cadavid, Radhi Datla, Suci Fitriany, Christine Haugen, Stephanie Mauro, Carissa Ikka Pardamean, Gauri Shanware, and Jyoti Singh for their invaluable contributions to the work presented in this paper.
Appendix A Specific Guidance for Paragraph and Sentence Creation
A.1 Overview
The Bouquet-source dataset needs to include 250 sentences in each of its source languages. This means that each linguist needs to create (i.e., create from scratch, not copy; see Section 2.4 above) 250 original sentences. These sentences must be:
  • Organized in logically structured paragraphs (see the Paragraphs section below)
  • Representative of the linguistic structures and features most frequently used in specific domains (see the Domains section below)
  • Representative of the most common register of language used in similar situations (see the Registers section below)
  • Accompanied by a gold-standard (i.e., best in class) human translation into English.
A.2 Paragraphs
The linguist will receive a template in the form of a spreadsheet, in which paragraph structures are designed and laid out. The template specifies the exact number of paragraphs and the exact number of sentences for each of the paragraphs. Each paragraph is given a unique paragraph ID (e.g., P01, P02, P15). Each sentence within each paragraph is also given a serial, non-unique ID (e.g., S1, S2, S3). The existing structure must not be modified in any way.
A.3 Domains
The template is divided into 8 domains:
  1. 1.
    How-to, written tutorials or instructions
  2. 2.
    Conversations (dialogues)
  3. 3.
    Narration (creative writing that doesn’t include dialogues)
  4. 4.
    Social media posts
  5. 5.
    Social media comments (reactive)
  6. 6.
    Other web content
  7. 7.
    Reflective piece
  8. 8.
    Miscellaneous (address to a nation, disaster response, etc.)
You must produce the set number of sentences for each of the domains; the domain / paragraph / sentence structure of the template cannot be changed.
A.4 Language Register Information
When creating sentences, make sure that the register of language being used is representative of the most expected and appropriate register for the situation. We understand that several registers could be possible; please use discretion when selecting a register, while making sure that the chosen register is among the most expected and appropriate. To help you make a determination, we define 3 main functional areas of language register:
  • Connectedness: What type of connection do language users who initiate the text have with other language users?
  • Preparedness: How much time do language users who initiate the text had or took to prepare the text?
  • Social differential: What is the relative social status of the language users who initiate the text towards other language users?
A.5 Linguistic Features
One of the main reasons for dividing the dataset into sections that correspond to domains is to attempt to cover as many registers and aspects of language as possible. For example, we know that:
  • Some pro-drop languages may drop the subject pronouns more often in some situations than in others.
  • Some case-marking languages may use some cases in specific situations but avoid them in others.
  • In English, lexical density increases when the level of formality increases.
  • Some languages use a specific past verb tense in storytelling, which stands out from other past verb tenses used in casual conversations or other situations.
  • Some languages use specific verb moods in some situations but avoid them in others.
A.6 Violating Content
While creating sentences, please be sure to avoid inserting violating content. Violating content is language that can fall under one (or more) of the below categories:
  • Toxicity
  • Illegal activities
  • Stereotypes and biases
A.7 Step-by-Step Description of Tasks
Please use the provided template to follow this step-by-step description.
Column A: Lang-ID This column should have the same 3-lowercase-letter code representing the source language of the sentences
being created followed by an underscore character ( _) and a 4-letter code representing the script.
Column B: Domain This is 1 of the 8 domains represented in the dataset (see Section 3.1).
Column C: Subdomain Please insert your description of the subdomain or topic.
Column D: P-ID This is the unique code identifying a paragraph (e.g., P01, P02, …, P58).
Column E: S-ID This is the non-unique code identifying the sequential place of the sentence within a paragraph.
Column F: Sentence In this cell, please type a sentence you created.
Column G: Translation into English After entering a sentence in your language in Column F, please provide a gold-standard human translation
in this cell.
Column H: S-Nchars This represents a count of the number of characters in the sentence.
Column I: S Comment_src_lang To help other linguists expand this dataset by translating your sentences into their own languages,
please add any comments that bring more context about the sentence.
Column J: S Comment_English Please provide an English translation of the comment your inserted in Column I.
Column K: Linguistic features Please list the register- or domain-specific linguistic features you tried to showcase in the sentence.
Column L: Connectedness Please use any of the options best describing the register area of Correctedness.
Column M: Preparedness Please use 1 of the options best describing the register area of Preparedness.
Column N: Social differential Please use any of the options best describing the register area of Social differential.
Column O: Formality Please indicate the level of formality best characterizing the sentence.
Column P: Relationship Please insert the intended relationship between the language users involved in the situation.
Column Q: Idea origin Please insert the name of the media type or platform that inspired the sentence.
Column R: P Comment_src_lang To help other linguists expand this dataset by translating your sentences into their own languages,
please add any comments that bring more context about the entire paragraph.
Column S: P Comment_English Please provide a translation into English for the comment you inserted in Column R.
Column T: P-Nchars This represents a count of the number of characters in the current paragraph.
Column U: Creator_Translator-ID Please insert your ID here, if it isn’t prepopulated.
A.8 Additional Guidance on Domain-Specific Content
Dialogues, especially those inserted in long creative writing (such as novels), often include the name of the speaker or a cue mark (e.g., — ), and sometimes quotation marks. When creating sentences for conversations, please feel free to invent names for speakers or to label speaker turns (e.g., A, B). Please place the names or speaker reference in markup tags, similarly to this: <Name:>or <A:>.
Emojis: There are emojis frequently in some social media and messaging domains. There should therefore be a representation, albeit very limited but the point isn’t to know how the system translates emojis because there are no real agreed ways to translate them across hundreds of languages.
Social media comments: Okay if the structure is flattened. Okay to have some tags but not absolutely necessary and no need to insert them everywhere.
Disfluencies in informal conversations?: If they are representative of conversations and they can be translated (i.e., there is some consensus on how to write them in the language — ah, oh, um).
Appendix B Contribution Statement
Pierre Andrews: Open initiative; Supervision/Leadership; core; tool deployment and design
Mikel Artetxe: Benchmarking; Supervision; support
Mariano Coria Meglioli: Dataset; core; statistical analysis
Marta R. Costa-jussa: Dataset, Benchmarking, Tooling Supervision/Leadership; core; corresponding author; research direction, paper editor
Joe Chuang: Dataset; core; source-bouquet creator
David Dale: Benchmarking; Supervision/Leadership; core; evaluation and experimental design
Cynthia Gao: Dataset; support
Jean Maillard: Open initiative; core, licensing
Alex Mourachko: Management; support
Christophe Ropers: Dataset and Open initiative; Supervision/Leadership; core; dataset design and creation, guidelines design and creation, paper editor
Safiyyah Saleem: Management; support
Eduardo Sánchez: Benchmarking; core; experiments deployment
Yiannis Tsiamas: Benchmarking; support
Arina Turkatenko: Dataset; core; dataset creation
Albert Ventayol-Boada: Dataset; core; dataset creation
Shireen Yates: Management; support

### v2

@@
 Abstract
-This paper presents BOUQuET, a multicentric and multi-register/domain dataset and benchmark, and its broader collaborative extension initiative.
-This dataset is handcrafted in non-English languages first, each of these source languages being represented among the 23 languages commonly used by half of the world’s population and therefore having the potential to serve as pivot languages that will enable more accurate translations.
-The dataset is specially designed to avoid contamination and be multicentric, so as to enforce representation of multilingual language features.
+BOUQuET is a multi-way, multicentric and multi-register/domain dataset and benchmark, and a broader collaborative initiative.
+This dataset is handcrafted in 8 non-English languages.
+Each of these source languages are representative of the most widely spoken ones and therefore they have the potential to serve as pivot languages that will enable more accurate translations.
+The dataset is multicentric to enforce representation of multilingual language features.
 In addition, the dataset goes beyond the sentence level, as it is organized in paragraphs of various lengths.
-Compared with related machine translation (MT) datasets, we show that BOUQuET has a broader representation of domains while simplifying the translation task for non-experts.
-Therefore, BOUQuET is specially suitable for the open initiative and call for translation participation that we are launching to extend it to a multi-way parallel corpus to any written language.
+Compared with related machine translation datasets, we show that BOUQuET has a broader representation of domains while simplifying the translation task for non-experts.
+Therefore, BOUQuET is specially suitable for crowd-source extension for which we are launching a call aiming at collecting a multi-way parallel corpus covering any written language.
+The dataset is freely available at https://huggingface.co/datasets/facebook/bouquet.
 1 Introduction
-Although multilingual large-language model evaluation benchmarks are only starting (Dac Lai et al., 2023), there is a rich research history in multilingual evaluation datasets for natural language processing; e.g., Sun and Duh (2020); Malmasi et al.
+Although multilingual large language model (LLM) evaluation benchmarks are only starting (Dac Lai et al., 2023), there is a rich research history in multilingual evaluation datasets for natural language processing; e.g., Sun and Duh (2020); Malmasi et al.
 (2022); Yu et al.
 (2022), with Machine Translation (MT) being the task with the highest investment in multilinguality (Kocmi et al., 2024).
-This is evident from the nearly 15-year history of the Conference on Machine Translation (formerly a workshop, WMT), which has established an international evaluation campaign (Kocmi et al., 2024).
-The campaign has compiled a comprehensive collection of parallel data evaluations covering a broad range of language pairs, domains, and tasks.
-Although some datasets offer multi-way parallelism across languages (e.g., the TED parallel corpus), [^1] https://hal.science/hal-04702210v1 the largest multi-way parallel corpus was only recently introduced with FLORES-101 (Goyal et al., 2022), later expanded to FLORES-200 (NLLB-Team et al., 2022) and to 2M-FLORES (Costa-jussà et al., 2024).
-However, even if great progress has been made, advances in multilingual applications for diverse languages are hindered by limited evaluations.
-Existing datasets and benchmarks fall short due to having an English-centric focus, a narrow register, compromised quality from automated construction and mining, limited language coverage, or a static nature, in addition to being prone to contamination.
-Similarly, in parallel with the previous progress, there have been several initiatives that called for data annotation in a collaborative and open way, such as the translation data collection initiative reported in Singh et al.
+This is evident from the nearly 20-year history of the Conference on Machine Translation (formerly a workshop, WMT), which has established an international evaluation campaign (Kocmi et al., 2024).
+The campaign has compiled a comprehensive collection of parallel corpus evaluations covering a broad range of language pairs, domains, tasks and recently, investing in a multi-way parallel dataset expanding in languages Deutsch et al.
 (2025).
-This precisely sets the stage for the introduction of a multilingual evaluation dataset and a benchmark, BOUQuET, which additionally combines community efforts.
-The purpose of this paper is two-fold.
-First, the paper details how we develop the Source-BOUQuET dataset, which is the necessary stepping stone towards an open initiative.
-Second, it presents how we design the open initiative itself, which aims to build the Full-BOUQuET dataset; i.e., Source-BOUQuET translated into any written language.
-1.1 Definitions
+However, the largest multi-way parallel evaluation dataset to date was introduced with FLORES-101 (Goyal et al., 2022), later expanded to FLORES-200 (NLLBTeam, 2024), FLORES+  [^1] https://oldi.org/ and to 2M-FLORES (Costa-jussà et al., 2024).
+These existing datasets and benchmarks fall short due to having an English-centric focus, a narrow selection of registers, compromised quality from automated construction and mining, limited language coverage, or a static nature, in addition to being prone to contamination Sainz et al.
+(2023).
+Similarly, in parallel with the previous progress, there have been several initiatives that called for data annotation in a collaborative and open way, such as the translation data collection initiative Singh et al.
+(2024).
+Recently, Wu et al.
+(2025) evaluate multilingual benchmarking and make a call for action for the need for accurate, contamination-free, challenging, practically relevant, linguistically diverse, and culturally authentic evaluations.
+This call and the urgent need of progressing in multilingual benchmarking set the stage for the introduction of a new multilingual multi-way parallel evaluation dataset and benchmark.
+BOUQuET, which additionally combines community efforts, relies on text written from scratch (contamination-free [^2] Note that BOUQuET is free from contamination in each initial state because it is originally created and not mined.
+However, from the moment we open-source certain splits, BOUQuET will risk to leak into training.
+Therefore, we keep one split hidden to avoid this.) by native speakers in 8 different major languages (linguistically diverse).
+Text includes a variety of 8 practical domains (practically relevant) that represent localised knowledge (culturally diverse).
+BOUQuET is aligned at the sentence and paragraph-level and it relies on a mixture of commissioned and openly collected human annotations to extend to any language.
+The organisation of the paper is as follows.
+First, the paper details how we develop the Source-BOUQuET dataset (Section 3), which is the necessary stepping stone towards an open initiative.
+Second, we benchmark BOUQuET for the 8 pivot languages plus English (Section 4).
+Finally, Section 5 presents how we design the open initiative itself, which aims to build the Full-BOUQuET dataset; i.e., Source-BOUQuET translated into any written language.
+At the time of submission of this paper, BOUQuET includes 55 multi-way parallel completed languages (Table 6).
+2 Definitions and background
+Definitions
 Before describing the Source-BOUQuET dataset’s characteristics and building methodology, we define our use of some frequently encountered terms that may cover a variety of meanings.
@@
 Register.
-We understand the term register as a functional variety of language that includes socio-semiotic properties, as expressed in Halliday and Matthiessen (2004), or more simply as a “contextual style”, as presented in Labov , pp.79–99).
+We understand the term register as a functional variety of language that includes socio-semiotic properties, as expressed in Halliday and Matthiessen (2004), or more simply as a “contextual style,” as presented in Labov (1991, pp.79–99).
 In that regard, a register is a specific variety of language used to best fit a specific communicative purpose in a specific situation.
+Background
+There is a large body of work in creating datasets for MT evaluation (e.g.
+WMT International Evaluation Campaigns Deutsch et al.
+(2025)).
+However, the vast majority are limited to a few languages.
+we next discuss the main efforts to build massively multilingual MT benchmarks and one representation of multi-domain dataset.
+FLORES+
+FLORES+ (Maillard et al., 2024) is the largest multilingual extension of FLORES-200 Goyal et al.
+(2022) and it covers the largest multi-way parallel dataset in terms of languages in 3 domains (Wikipedia, News, Travel guides).
+Even if FLORES+ has paragraph information, the translation has been done at the level of sentence without showing context to the annotators.
+NTREX-128
+Similarly to FLORES+ NTREX-128 covers a multi-way parallel dataset but for 128 languages.
+Unlike FLORES-200, translators had the full context of the document available when translating sentences, but the authors did not know if (or to what extent) they used this information Federmann et al.
+(2022).
+NLLB-MD
+was motivated to complement FLORES-200 in terms of domains in the context of the NLLB NLLBTeam (2024) project.
+It covers chat, news and health domains in 6 languages and it includes a much larger number of sentences.
+All these datasets are English-localised and English-centric, meaning that all languages have been translated from the original source English.
+They cover limited amount of domains (a maximum of 4) and do not differentiate among registers.
 3 Dataset: Source-BOUQuET
-In this section, we describe the creation criteria that have follow to design Source-BOUQuET, as well as the languages it includes.
+In this section, we describe the creation criteria that have been followed to design Source-BOUQuET, as well as the languages it includes.
 3.1 Main characteristics
-As described in greater detail next, the Source-BOUQuET dataset is mainly characterized by its non-English-centric focus, its diverse registers and domain (which is complementary to FLORES-200), its manual and original composition, and its built-in dynamic extensibility.
+As described in greater detail next, the Source-BOUQuET dataset is mainly characterized by its non-English-centric focus, its diverse range of registers and domains (which are complementary to FLORES-200), its manual and original composition, and its built-in dynamic extensibility.
+Table 1 provides a comparison of several relevant statistics from BOUQuET and the closest related datasets covered in the previous section.
 Non-English-centric focus.
-Source-BOUQuET is handcrafted by proficient speakers of French, German, Hindi, Indonesian, Mandarin Chinese, Russian, and Spanish.
-Each of those languages provides the same number of sentences to the final dataset.
-The languages for Source-BOUQuET (see Table 2 in Section 2.3) are all part of the top 20 languages in the world in terms user population, as listed in Eberhard et al.
+Source-BOUQuET is handcrafted by proficient speakers of Egyptian Arabic and MSA, French, German, Hindi, Indonesian, Mandarin Chinese, Russian, and Spanish.
+Each of these languages contributes the same number of sentences to the final dataset.
+The languages for Source-BOUQuET (see Table 3 in Section 3.3) are all part of the top 20 languages in the world in terms of user population, as listed in Eberhard et al.
 (2024).
-In addition, they are also used by a large number of both native and non-native speakers, which makes them good candidates for what we refer to as pivot languages; i.e., higher-resource languages that can facilitate—as source languages—the translation of datasets into lower-resource languages.
+In addition, they are also used by a large number of non-native speakers, which makes them good candidates for what we refer to as pivot languages; i.e., higher-resource languages that can facilitate—as source languages—the translation of datasets into lower-resource languages.
 English is often used as such a pivot language, since numerous people have a high degree of proficiency in English as a second language.
@@
 However, the relationship between registers and domains is not one to one.
-For example, if we take a domain such as TV news, we can identify at least 3 registers:
-  • The register used by the news anchor, which is represented by fully scripted language that is read from a teleprompter with a very specific and unnatural form of diction (hypercorrect enunciation, unnatural intonation, homogeneous pace).
-  • The register produced by communication specialists (i.e., people who have been trained to be spokespersons or surrogates).
+See the register and domain correspondence in Figure 5 (Appendix B).
+For example, if we take a domain such as TV news, we can identify at least 3 registers: (1) the register used by the news anchor, which is represented by fully scripted language that is read from a teleprompter with a very specific and unnatural form of diction (e.g., hypercorrect enunciation, unnatural intonation, homogeneous pace); (2) The register produced by communication specialists (i.e., people who have been trained to be spokespersons or surrogates).
 The points they make have been scripted and rehearsed to the point of being known by heart.
-Communication specialists also learn rehearsed patterns of language to make sure they do not get surprised by a question and can redirect the conversation towards the points they need to make instead of answering unexpected questions.
-All the language they use has been written and rehearsed.
-It sounds spontaneous but it is not structured like informal language.
-  • The register represented in person-in-the-street segments, which is more informal and spontaneous (possibly colloquial).
-The example above is taken from a domain where both speech and writing are used but the situation is not significantly different in the written modality only.
+It sounds spontaneous but it is not structured like informal language; (3) the register represented in person-in-the-street segments, which is more informal and spontaneous (possibly colloquial).
+This example is taken from a domain where both speech and writing are used but the situation is not significantly different in the written modality only.
 Language users all commonly shift between registers, which is typically referred to as style-shifting.
 Style-shifting (i.e., register-shifting) occurs within domains; so the domain itself is not a fool-proof way of getting a specific register.
-Although the norms of the domain can impose the degree of formality and of lexical specialization, it is often the register (which derives from the communicative purpose), not the domain, that determines many aspects of linguistic structure (lexical density, pronoun use, syntax, etc.).
-The registers we cover is Source-BOUQuET are given in Figure 1.
+Although the norms of the domain can impose the degree of formality and of lexical specialization, it is often the register (which derives from the communicative purpose), not the domain, that determines many aspects of linguistic structure (e.g., lexical density, pronoun use, syntax, etc.).
 Manual construction and original composition (not crawled) with accurate revisions
-To develop Source-BOUQuET, we set a variety of linguistic criteria that need to be covered.
-Guidelines are then shared with linguists who manually craft sentences covering examples of these linguistic criteria and compose paragraphs of various lengths.
-The main strategies for open collaboration are to design contribution guidelines and build the annotation tool that enables to freely collect translations in any language.
-For this, BOUQuET is initially composed on 8 major languages plus the English translation and it is shared in a repository that allows language communities to easily add a new language even if they do not speak English, e.g.
-a native speaker of Quechua can translate from Spanish.
+To develop Source-BOUQuET, we set a variety of linguistic criteria that need to be covered, including both unmarked and marked structures (e.g., expected and unexpected number agreement between subject and verb).
+Guidelines are then shared with linguists who manually craft sentences covering examples of these linguistic criteria and compose paragraphs ranging from 3 to 6 sentences in length.
+These paragraphs are then manually translated across all pivot languages.
+The main strategies for open collaboration are to design contribution guidelines and build an annotation tool that enables the free collection of translations in any language.
+BOUQuET is shared in a repository that allows language community to easily add a new language by translating it from one of the 8 pivot languages or the English translation.
 This repository contains detailed guidelines on how to do it.
-BOUQuET’s innovative approach ensures universal language accessibility.
+BOUQuET’s innovative approach ensures widespread language accessibility.
 This open collaborative initiative will enrich BOUQuET with the following characteristics.
@@
 Since BOUQuET includes the community, it can continuously evolve by constantly engaging it.
+Dataset Split ##\##Parag. ##\##Sent Avg. Wrd. Parag/Sent Reg. Dom. Lang. Dyn.
+FLORES+ Dev $\times$× 997 Wikipedia, News, Travel guides 220
+Devtest 1,012 25 $\times$×
+Eval 992
+NTREX-128 Test 123 1,997 389/24 $\times$× News 128 $\times$×
+Dev 6,000
+NLLB-MD Devtest $\times$× 1,310 25 $\times$× Chat, News, Health 6 $\times$×
+Eval 1,500
+BOUQuET Dev 120 504 Fiction, Conversation, Social media 55+ [^3] See Appendix D for language coverage details
+Devtest 200 864 55/15 posts/comments, Tutorials, Website,
+Eval 144 628 Reflection pieces, Miscellaneous
+Table 1: Main statistics from MT evaluation datasets including BOUQuET: number of sentences, number of paragraphs, average word per paragraph (or sentence), register information, domains, languages, dynamism.
 3.2 Creation criteria
-For the design of the creation guidelines, which are reported in detail in Appendix A, we prepared a list of linguistic coverage requirements as well as some statistical information.
+For the design of the creation guidelines, detailed in Appendix A, we prepared a list of linguistic coverage requirements along with some statistical information.
 Linguistic coverage requirements.
 In order for BOUQuET to be representative of various linguistic phenomena, linguistic coverage requirements are defined (as listed in Table 2), which are to be included in sentences that form paragraphs.
-This continuity will include variation in the number of sentences per paragraph and in the lengths of sentences themselves.
+Sentences are assigned a unique identifier that combines a unique paragraph ID number with a serial sentence number.
+Thus, paragraphs can be retrieved by concatenating sentences that share the same paragraph ID.
 Phenomena
@@
 Different emphasis or topicalization
-Different sentence structures (affirmation, interrogation, negation, subordination, coordination)
+Different sentence structures (affirmation,
+interrogation, negation, subordination, coordination)
 Different verb moods, tenses, and aspects
@@
 Variety of domains.
-Source-BOUQuET is intended to cover 8 domains: narration (as in fiction writing), dialog, social media posts, social media comments, how-to, miscellaneous website content (excluding social media or news), opinion pieces, miscellaneous other (such as written speeches or signage).
+Source-BOUQuET is intended to cover 8 domains: narration (as in fiction writing), dialog, social media posts, social media comments, how-to manuals and instructions, miscellaneous website content (excluding social media or news), opinion pieces, and other miscellaneous (such as written speeches or signage).
 The choice of these domains optimizes for variety and popular usefulness.
@@
 Source-BOUQuET is built with register variety in mind, differently from FLORES-200, which covers a few different domains but remains largely within similar registers.
-We characterize the registers through 3 main aspects (connectedness, preparedness, and social differential), and a variety of options within these aspects (see Figure 1 for detail).
+We characterize the registers through 3 main features (connectedness, preparedness, and social differential).
 Connectedness attempts to describe the type of interaction typically available in a given domain.
 Preparedness aims to gauge how much time is typically used to produce or edit language content.
-Social differential tries to describe various social situations in which language content is typically produced in each of the domains.
-Options are not mutually exclusive at the level of the domain but become differentiated at the level of a single sentence.
+Social differential describes the relationship between the interlocutors involved in a given social situation (e.g., writer and reader, characters in a dialog, etc.).
+Each individual domain can present different combinations of features but become differentiated at the level of the sentence.
+There are a variety of feature combinations, which are mentioned in Figure 5 and defined in Appendix B.
 By including new registers and domains, the new dataset is likely to be more generalizable to different contexts and applications.
-Figure 1: Covered register characteristics.
-Statistical guidance.
-In order to appropriately cover linguistic requirements and adequately represent the registers and domains, we performed a statistical analysis to understand each of the linguistic characteristics.
-In particular, our analysis covers most domains that we are including in Source-BOUQuET by using diverse public datasets: narration (Books3, gutenberg (Gerlach and Font-Clos, 2018)); Social media posts (Reddit (Baumgartner et al., 2020)); Social media comments (Wikipedia comments [^2] https://www.kaggle.com/competitions/jigsaw-multilingual-toxic-comment-classification); Conversations / Dialogues (dialogsum (Chen et al., 2021), Open Orca (Lian et al., 2023)); Tutorials/how-to articles (how-to Wikipedia-lingua  [^3] https://huggingface.co/datasets/GEM/wiki_lingua); Website content (c4 (Raffel et al., 2020)); News / Reflection pieces (cnndailymail (Nallapati et al., 2016), xsum (Narayan et al., 2018)) and Miscellaneous (Wikipedia).
+Statistical guidance for domain representation.
+In order to appropriately cover linguistic requirements and adequately represent domains, we performed a statistical analysis to understand the linguistic characteristics of each domain before creating BOUQuET.
+In particular, our analysis covers most domains that we are including in Source-BOUQuET by using diverse public datasets: narration (Books3, Gutenberg library (Gerlach and Font-Clos, 2018)); Social media posts (Reddit (Baumgartner et al., 2020)); Social media comments (Wikipedia comments [^4] https://www.kaggle.com/competitions/jigsaw-multilingual-toxic-comment-classification); Conversations / Dialogues (dialogsum (Chen et al., 2021), Open Orca (Lian et al., 2023)); Tutorials/how-to articles (how-to Wikipedia-lingua  [^5] https://huggingface.co/datasets/GEM/wiki_lingua); Website content (C4 (Raffel et al., 2020)); News / Reflection pieces (CNN-DailyMail (Nallapati et al., 2016), XSum (Narayan et al., 2018)) and Miscellaneous (Wikipedia).
 Note that we collect information from public data that do not always accurately match our categories but constitute a proxy.
-For each of these domains, we have analyzed dimensionality: characters per token; tokens per sentence and sentences per paragraph.
-Figure 2: (Left) Tokens per sentence and (middle) sentences per paragraph (Right) CEFR per domain.
+For each of these domains, we have analyzed dimensionality: characters per token; tokens per sentence and sentences per paragraph; and linguistic complexity with CEFR levels [^6] https://www.coe.int/en/web/common-european-framework-reference-languages/.
+Figure 1: (Top) Tokens per sentence and (Middle) sentences per paragraph (Bottom) CEFR per dataset representative of BOUQuET domains.
 Regarding tokens per sentence (Figure 1 left), we can see correlations between different domains, and clear differences in length, especially in dialogs which tend to be much shorter.
-Regarding sentences per paragraph (Figure 2 middle), we can find correlation between different datasets representing the same domain, where fiction writing paragraphs tend to be much longer (averaging 5 but reaching up to 20 ), dialogs and news articles are much shorter (barely reaching 3-4 sentences in a paragraph), and the rest of the categories are somewhere in between (normally staying between 1-5 but reaching up to 10 in some cases).
-Furthermore, we have checked complexity: distribution of CEFR level [^4] https://rm.coe.int/common-european-framework-of-reference-for-languages-learning-teaching/16809ea0d4 as a proxy.
-Including here the % of C2 scores at sentence level for each dataset.
-This was labeled by a sonar-based model (Duquenne et al., 2023) trained on CEFR-SP data CEFR Text Classification which greatly outperformed Llama-3.1 (Touvron et al., 2023).
-See Figure 2 right.
+Regarding sentences per paragraph (Figure 1 middle), we can find a correlation between different datasets representing the same domain, where fiction writing paragraphs tend to be much longer (averaging 5 but reaching up to 20 ), dialogs and news articles are much shorter (barely reaching 3-4 sentences in a paragraph), and the rest of the categories are somewhere in between (normally staying between 1-5 but reaching up to 10 in some cases).
+To guide BOUQuET creators on the linguistic complexity required for each domain, we have assessed complexity using the distribution of CEFR levels as a proxy.
+This includes the % of C2 scores at the sentence level for each dataset.
+These scores were labeled by a SONAR-based model (Duquenne et al., 2023) trained on CEFR-SP data for CEFR Text Classification, which significantly outperformed Llama-3 (Touvron et al., 2023).
+See right side of Figure 1.
 Wikipedia seems to be the only dataset with a more considerable share of C2 sentences, with some others like dialogues having no samples scored as such.
-Annotations.
-Each entry of Source-BOUQuET includes the source text (in one of the 8 pivot languages from Table 2) and its translation into English, domain information, and contextual information for better translation accuracy.
+Annotations and Quality Checks
+Each entry of Source-BOUQuET includes the source text (in one of the 8 pivot languages of Table 3) and its translation into English, domain information and contextual information for better translation accuracy.
+To double-check that Source-BOUQuET does not contain repeated sentences, we explored the similarity across English sentences.
+For each English sentence, we computed SONAR embeddings Duquenne et al.
+(2023) and we computed the cosine distance on the vectors.
+There were only 14 sentences with a cosine distance below 0.3.
+These sentences are reported in Appendix E.
 3.3 Languages
-BOUQuET aims to be as multicentric as possible, in contrast with most existing datasets like FLORES-200 that are English-centric.
+As mentioned earlier, BOUQuET aims to be multicentric and localized , in contrast to most existing datasets that are English-centric.
 The motivation is mainly to be representative of linguistic phenomena.
-To this effect, it is created in 7 non-English languages.
+To this effect, it is created in 8 non-English languages (Table 3).
 Each language contributes with a similar number of sentences along with their English equivalents given by the sentence creators themselves.
-See languages in Table 2.
-ISO 639-3 ISO 15924 Language Family Subgroup1
-cmn Hans Mandarin Chinese Sino-Tibetan Sinitic
-deu Latn German Indo-European Germanic
+3.4 Multi-way extension to Source-BOUQuET languages
+Details
+Source-BOUQuET creators composed 250 sentences for each of the 8 pivot languages plus the corresponding English translation.
+The remaining 1,750 sentences for each pivot language are translated from English.
+The final Source-BOUQuET is composed of 2,000 sentences in 9 languages (8 pivot languages plus their translations into English).
+Quality checks
+Since multi-way parallel data is created from English, we manually checked that translations did not lose the linguistic information when translating from English.
+While translating BOUQuET, we had to make sure that the contextual information which was applicable to the whole paragraph was taken into consideration by the translators.
+To ensure this, we used a number of following QA strategies reported in Appendix C.
+Additional contextual information
+The multi-centric nature of BOUQuET is also a reminder that English is not morphologically rich (e.g., it doesn’t mark grammatical gender agreement between nouns, adjectives, and verbs) and displays relatively little information about formality in its written form (e.g., it uses only one second-person singular pronoun, regardless of who is addressing whom).
+As such, English isn’t an ideal source language for translation purposes unless translators can be provided with additional contextual information.
+The BOUQuET dataset includes such additional information; for example, the grammatical gender of the first and second person (when this isn’t obvious) or the linguistic markedness of some words or phrases (e.g., literary or archaic verb tenses, use of slang, infrequently used level of formality).
+ISO ISO Language Family Subgroup1
+6393 15924
+arb Arab Modern Afro-Asiatic West Semitic
+Standard Arabic
+cmn Hans Mandarin Sino-Tibetan Sinitic
+Chinese
+deu Latn German Indo-European West Germanic
 fra Latn French Indo-European Italic
 hin Deva Hindi Indo-European Indo-Aryan
-ind Latn Indonesian Austronesian Malayo-Polynesian
+ind Latn Indonesian Austronesian Malayic
 rus Cyrl Russian Indo-European Balto-Slavic
@@
 Table 3: Source-BOUQuET Languages
+3.5 Overall Statistics
+In total, BOUQuET contains 2,000 sentences.
+These sentences are split by making a stratified paragraph-level selection among source languages and domains into development, test and evaluation sets.
+Initially, the evaluation set (632/144 sentences/paragraphs) is intended to be kept hidden.
+Figure 2 shows the representation of registers (top) and domains (bottom) in the non-hidden splits.
+Labels for each of the combinations of register options are created by concatenating the lowercase letters used as unique identifiers (see details of these register options in the Appendix B).
+For example, a register characterized as impersonal (in connectedness), composed (in preparedness), and equal-assumed (in social differential) is labeled ica.
+Figure 2: Registers (top) and domains (bottom) representations in development and test partitions.
+The results in the following section are presented with the test split of 864/200 sentences/paragraphs.
 4 Benchmark
-Statistics
-In total, BOUQuET currently contains 1,750 sentences.
-These sentences are split by doing a stratified selection among source languages and domains into development and test sets with an equivalent number of sentences.
-A portion of the development set (522) is intended to initially be kept hidden.
-Results in the following sections are presented with the test split of 889 sentences.
-Table 3 provides a comparison of several relevant statistics from BOUQuET and the closest related datasets including FLORES-200, NTREX-128 (Federmann et al., 2022) and NLLB-MD (NLLB-Team et al., 2022).
-Dataset Sent. Parag. Reg. Dom. Lang. Dyn.
-FLORES-200 3,003 () $\times$× Wikipedia, News, Travel guides 200 $\times$×
-NTREX-128 1,997 $\times$× News 128 $\times$×
-NLLB-MD 9,000 $\times$× $\times$× Chat, News, Health 6 $\times$×
-BOUQuET 1,750 Fiction, Conversation, Social media posts/comments 8+
-Tutorials, Website, Reflection pieces, Miscellaneous
-Table 3: Main statistics from MT evaluation datasets including BOUQuET: number of sentences, paragraph-information, registers, domains, languages, dynamism.
+We benchmark BOUQuET in two dimensions: domain representation and machine translation.
+The former quantifies how representative BOUQuET is of public datasets of multiple domains compared to other evaluation datasets.
+The latter addresses how several MT systems are ranked with BOUQuET compared to other evaluation datasets.
 Domain representation
 The performance of the model in a new or unseen dataset depends on the similarity between the dataset that was used to fit the model and the new dataset.
-In order to more broadly evaluate an MT model, we are interested in covering a different set of domains.
-We compare the domain coverage of BOUQuET with that of FLORES-200, NTREX-128 (Federmann et al., 2022) and NLLB-MD (Stolte et al., 2024) .
-To do this comparison, we represent data (multi-domains public datasets from Figure 2 in grey, each alternative dataset FLORES-200, NTREX-128 and NLLB-MD and BOUQuET) with SONAR (Duquenne et al., 2023).
-From SONAR vectors, we do a PCA-dimensionality reduction, see Figure 3.
-All domains from Figure 2 are represented in grey.
-Then, Figure 3 from left to right compares BOUQuET against FLORES-200, NTREX-128 and NLLB-MD, respectively.
+We compare the domain coverage of BOUQuET with that of FLORES+, NTREX-128 and NLLB-MD.
+To do this comparison, we take a random sample of 2,000 sentences (which seems to be a sufficiently large sample of the embedding space for score stability) from each of the domain datasets from Figure 1; as well as 2000 from each alternative dataset FLORES+, NTREX-128, NLLB-MD,and BOUQuET.
+We create vector representations of each sentence in previous datasets with SONAR (Duquenne et al., 2023).
+From SONAR vectors, we do a PCA-dimensionality reduction, fitted upon the combined multi-domain set, see Figure 6 (Appendix F).
+Public domains from Figure 1 are represented in grey; alternatives evaluation datasets are represented in blue and BOUQuET is represented in red.
+Figure 6, from top to down, compares BOUQuET against FLORES-200, NTREX-128, NLLB-MD, respectively.
 We qualitatively observe that BOUQuET covers a wider range of domains.
-Additionally, to quantify this coverage, we measure the overlap between each dataset with each of the domains using two metrics, the Wasserstein distance (implemented with POT library [^5] https://pythonot.github.io/) and the Silhouette coefficient (implemented with sklearn [^6] https://scikit-learn.org/stable/).
-The Wasserstein Distance (WD), also known as the Earth Mover’s Distance (EMD), is a metric that measures the "effort" required to transform one probability distribution into another.
+Additionally, to quantify this coverage, we measure the overlap between each dataset with each of the domains using the Wasserstein distance (implemented with the POT library [^7] https://pythonot.github.io/).
+The Wasserstein Distance (WD), also known as the Earth Mover’s Distance (EMD), is a metric that measures the ”effort” required to transform one probability distribution into another.
 Lower results indicate a higher similarity between clusters.
-The lowest consistent results for all domains are obtained with the data set BOUQuET.
-Alternatively, the Silhouette Score ($S$)  [^7] https://en.wikipedia.org/wiki/Silhouette_(clustering) measures how well each data point fits within its assigned cluster compared to other clusters, it ranges between -1 and 1 and the higher the coefficient, the more coherent the clusters are.
-Again, BOUQuET shows a lower score, consistent across domains, meaning closer representation across domains.
-Figure 3: Domain representation and overlap with datasets (FLORES-200, NTREX-128, NLLB-MD)
-Narration Website Opinion Dialog Social media Social media How-to Misc.
-comments posts
-WD SC WD SC WD SC WD SC WD SC WD SC WD SC WD SC
-FLORES-200 0.32 0.05 0.32 0.03 0.32 0.02 0.31 0.11 0.32 0.09 0.31 0.08 0.31 0.06 0.33 0.03
-NTREX-128 0.32 0.04 0.31 0.03 0.32 0.01 0.31 0.09 0.31 0.07 0.31 0.6 0.31 0.06 0.33 0.03
-NLLB-MD 0.31 0.04 0.31 0.02 0.31 0.00 0.30 0.08 0.31 0.06 0.30 0.05 0.30 0.05 0.32 0.02
-Bouquet 0.29 -0.02 0.29 -0.02 0.30 -0.03 0.28 0.02 0.29 0.01 0.28 0.00 0.28 0.00 0.30 -0.01
-Table 4: Wasserstein Distance (WD) and Silhouette Score ($S$) for each domain and dataset. Lower WD and $S$ indicate better representation of the domain. Best results in bold.
-MT Benchmarking
+This distance is run on the full 1024-dimensional SONAR embedding vectors, without applying any kind of dimensionality reduction (PCA).
+Some domain sets and evaluation sets were several orders of magnitude larger than each other.
+Sampling all down to the same size (2,000) makes the metric computable in a reasonable amount of time and removes any sensitivity to class imbalance in the distribution distance metric.
+Figure 3 shows that the lowest consistent results are obtained for all domains with BOUQuET.
+Figure 3: Wasserstein Distance (WD) for each domain and dataset. Lower WD indicate better representation of the domain.
+Machine Translation
 To help the reader understand why the dataset is useful, we present preliminary results to demonstrate its use for its intended purpose: MT benchmarking.
-We evaluated on translation systems (GPT-4o (Wu et al., 2024), Llama-3.1 (Llama-3.3-8B)(Touvron et al., 2023), and NLLB (NLLB-3.3B) (NLLB-Team et al., 2022)) using two metrics: Comet (CometKiwi-da-xl, range 0-1 and $\uparrow$ better) (Chimoto and Bassett, 2022) and MetricX (MetricX-24, range 0-25 and $\downarrow$ better) (Juraska et al., 2024), following WMT 2024 (Kocmi et al., 2024).
-See results in Table 5.
-We observe that BOUQuET scores consistently higher than NLLB-MD on average and for all models, suggesting that BOUQuET is easier to translate.
+We evaluate 14 translation systems: Llama-3 (Llama3.1-8B, Llama3.2-3B, Llama3.3-70B)(Touvron et al., 2023), Tower (TowerInstruct-7B-v0.2) (Rei et al., 2024), Aya (Aya101-13B, Aya-Expanse-8B) (Dang et al., 2024), Babel (Babel-9B-Chat) (Zhao et al., 2025), Cohere (CohereLabs-command-r7b-12-2024), Eurollm (EuroLLM-9B-Instruct) (Martins et al., 2024), MADLAD (MADLAD-3B-MT and MADLAD-10B-MT) (Kudugunta et al., 2023), Mistral (Mistral-7B-Instruct-v0.3) [^8] https://docs.mistral.ai/getting-started/models/models_overview/, Qwen (Qwen2.5-7B-Instruct) (Bai et al., 2023) and NLLB (NLLB-3.3B) (NLLBTeam, 2024).
+We select the models as ones with open weights, focusing primarily on moderate sizes (about 10B) and variety of architectures.
+Following the official evaluation metrics of WMT 2024 (Kocmi et al., 2024), we use two automatic metrics: CometKiwi (CometKiwi-da-xl, range 0-1 and $\uparrow$ better, COM) (Chimoto and Bassett, 2022) and MetricX (MetricX-24-hybrid-xl-v2p6, range 0-25 and $\downarrow$ better, MetX) (Juraska et al., 2024).
+We include in the benchmarking datasets that cover Source-BOUQuET languages (FLORES+ and NTREX-128).
+Table 4 shows that BOUQuET scores consistently higher than other datasets on average, suggesting that BOUQuET is easier to translate.
 This is an advantage for the open initiative, since the complexity of current MT test sets makes it harder to ask the community to participate in translations as it requires a high-level of expertise.
-Rankings across models nor languages is not preserved, which hints that BOUQuET may be posing different challenges to the models.
+Rankings across models and datasets is not preserved, which hints that all datasets may be posing different challenges to the models.
+Rankings is computed as counting when a system is similar in the same position according to CometKiwi.
+This ranking and Pearson correlation on the CometKiwi is dissimilar for datasets evaluated at the sentence-level, with BOUQuET being the most different.
+This difference is enlarged when evaluating at the paragraph-level where number of swaps increases and, coherently, Pearson correlation decreases, meaning that datasets pose different challenges to models.
 We need to further investigate which linguistic challenges BOUQuET is adding.
-On this ranking, if we look at average BOUQuET, LlaMa is the worst system according to Comet but the best system according to Metricx.
-This does not happen in NLLB-MD, where the ranking is consistent across metrics.
-If we look per language, ranking across metrics is consistent for 4 out of 7 languages for NLLB-MD and for 1 out of 7 languages for BOUQuET.
-This raises a concern and may suggest that MT metrics themselves do not generalize well to the broader domains/registers in BOUQuET.
-However, this requires further investigation and we want to specifically explore different alternatives to use BOUQuET such as exploring the performance by domain and register.
-Note that we present only preliminary experiments, see Section 5 for more details on ongoing experiments.
-M cmn deu fra hin ind rus spa AVG
-COM($\uparrow$) MetX($\downarrow$) COM MetX COM MetX COM MetX COM MetX COM MetX COM MetX COM MetX
-NLLB-MD
-GPT4o 0.76 14.34 0.77 6.80 0.76 13.86 0.68 6.63 0.79 16.16 0.78 3.96 0.80 13.99 0.76 10.81
-Llama 0.71 14.38 0.71 6.96 0.72 14.00 0.57 7.07 0.75 16.21 0.71 4.00 0.75 14.11 0.70 10.96
-NLLB 0.61 14.80 0.70 7.01 0.70 13.98 0.64 6.72 0.73 16.20 0.72 4.43 0.76 14.06 0.69 11.03
-BOUQuET
-GPT4o 0.79 1.99 0.80 2.79 0.78 5.57 0.76 3.37 0.84 4.07 0.81 3.60 0.83 4.74 0.80 3.73
-Llama 0.75 2.06 0.76 2.07 0.74 4.68 0.72 3.19 0.83 3.54 0.75 3.26 0.80 4.07 0.76 3.27
-NLLB 0.67 2.19 0.79 2.44 0.77 5.30 0.74 3.49 0.82 3.94 0.78 3.40 0.67 4.42 0.77 3.60
-Table 5: Results on Quality Estimation from English to 7 Source-BOUQuET languages on BOUQuET evaluation dataset .
-4 Open initiative
+However, best two systems are consistent across datasets and level of evaluation (sentence and pargraphs) being those the largest model (Llama3.3-70B) and Aya-e-8B.
+NLLB-3.3B has a higher variation between being evaluated at the sentence or paragraph-level, which makes sense since it is the only one trained with sentence-level data.
+Model Bouquet Flores Ntrex BouquetP NtrexP
+COM MetX COM MetX COM MetX COM MetX COM MetX
+NLLB-3B 0.68 2.1 0.66 2.56 0.65 2.97 0.59 3.71 0.29 14.1
+aya101-13B 0.67 2.02 0.63 2.65 0.63 3.14 0.58 3.29 0.24 13.17
+aya-e-8B 0.69 1.75 0.65 2.9 0.67 2.45 0.64 2.42 0.34 8.7
+babel-9B 0.67 2.33 0.65 2.66 0.63 3.36 0.61 3.4 0.32 10.39
+cohere-7B 0.67 2.15 0.65 2.89 0.64 3.01 0.61 3.2 0.32 9.61
+eurollm-9B 0.67 2.33 0.65 2.89 0.61 3.64 0.61 3.64 0.31 10.08
+madlad-10B 0.63 2.74 0.64 2.72 0.63 3.35 0.41 6.76 0.15 15.99
+madlad-3B 0.63 2.85 0.63 2.94 0.61 3.67 0.37 6.71 0.49 5.29
+mistral-7B 0.54 4.29 0.51 5.69 0.49 6.12 0.49 6.64 0.24 10.96
+qwen-7B 0.59 3.25 0.6 3.75 0.59 4.21 0.57 4.5 0.52 4.93
+Llama3.1-8B 0.66 2.36 0.64 2.82 0.63 3.27 0.6 3.33 0.32 10.17
+Llama3.2-3B 0.59 3.59 0.57 4.34 0.55 4.89 0.52 5.52 0.27 12.67
+Llama3.3-70B 0.7 1.85 0.68 2.21 0.67 2.59 0.63 2.72 0.35 9.76
+Tower-7B 0.58 3.69 0.56 4.19 0.56 4.35 0.49 5.65 0.28 12.22
+Bouquet-Flores Flores-Ntrex Ntrex-Bouquet BouquetP-NtrexP
+Swaps 3 4 7 11
+Pearson Cor. 0.95 0.99 0.95 0.92
+Table 4: Averaged Results $X$-to-$X$ 9 Source-BOUQuET (8 pivot plus English) languages for BOUQuET, FLORES+, NTREX-128 at the level of sentence 2 columns on the left and at the level of paragraph 3 columns on the right.
+Number of ranking swaps (a system not being in the same position according to CometKiwi) from each dataset compared to the other two (in similar sentence or paragraph-level) and Pearson correlation indicate that while datasets report similar results at sentence-level, being BOUQuET the most different, it is not the case for paragraph-level where the ranking of systems varies by a larger amount.
+Figure 4 shows results of the 3 best systems averaged across language directions, evaluated at the sentence-level, per domains.
+Worse performing domains are comments, conversations, how-to and narration.
+Best performing domains are web and other miscellaneous, reflection and social posts.
+Appendix F reports more detailed results on BOUQuET per language and domains.
+Figure 4: Best performing models and their results in each of the BOUQuET domains .
+5 Beyond commissioning translations: Open initiative
 Source-BOUQuET is intended to be translated into any written language.
+For this, we have commissioned an initial set of priority languages covering a variety of high and low-resource languages representing different geographical regions, linguistic families and scripts.
+See the list of languages currently covered by BOUQuET in Appendix D.
+However, it would be challenging to achieve our language coverage target to any language.
 This ambition can only be achieved with the support of the community.
@@
 The purpose of this open initiative is to collect translations from Source-BOUQuET.
-To collect these annotations, we have chosen the ARGILLA tool [^8] https://argilla.io/ which is a collaboration platform for building datasets, designed for use by AI engineers and domain experts.
-Together with setting the Source-BOUQuET in this tool, we have designed annotation guidelines which very much resemble those from FLORES-200 (NLLB-Team et al., 2022).
-This open initiative is available in HuggingFace https://huggingface.co/spaces/facebook/bouquet and it is starting as part of the Meta Language Technology Partner Program [^9] https://docs.google.com/forms/d/e/1FAIpQLSdzcRdtkQCuTrXw727DgJgWbOPKDj5v0bArgGfQUTT6sEopFw/viewform publicly launched in AI Summit [^10] https://www.elysee.fr/en/sommet-pour-l-action-sur-l-ia.
+To collect these annotations, we have set a tool to collect annotations.
+Together with setting Source-BOUQuET in this tool, we use the annotation guidelines from Section 3.4 which very much resemble those from FLORES-200 (NLLBTeam, 2024) and which are available in the 9 BOUQuET languages.
+One of the advantages is that annotators can choose the source language from among one of the Source-BOUQuET languages.
+These languages have been chosen to cover a wide range of speakers, facilitating the task of annotation instead of depending on English bilingual speakers.
+This open initiative is available in https://bouquet.metademolab.com/.
 6 Conclusions and Next Steps
 In this paper, we have presented the Source-BOUQuET dataset and the attached open initiative.
-This dataset is available in 8 languages.
-We have shown that consistent BOUQuET gains in domain diversity by in two different metrics while keeping complexity lower than its competitors.
-The latter is specially relevant to simplify the translation for non-experts that may join the open initiative.
-This paper and Source-BOUQuET is only a fraction of the ambitious project that we are starting by launching this open call for community efforts.
-Please join us in making Universal Quality Evaluation in Translation available in any language!
-Limitations
-BOUQuET is work in progress; the BOUQuET dataset is still limited in number of languages and translations.
-The benchmarking is also limited in several axes (dataset comparison, models and metrics evaluated).
-However, we prioritized starting the BOUQuET open initiative in coordination with the Meta Language Technology Partner Program launch so that we could start raising the collaboration interest.
-As OmnilingualMT team, we continue actively investing on it, see next steps that we are planning to release soon.
-Coming Soon
-BOUQuET is actively evolving and we are currently working on making the following progress:
-  • Adding 250 sentences in Source-BOUQuET in Egyptian Arabic.
-  • Complete Source-BOUQuET in all 7 languages from Table 2 plus Egyptian Arabic.
-  • The creation guidelines plus translation guidelines in all 8 pivot languages.
-  • Designing quality control for each of the contributions and adding new languages to the incremental releases of BOUQuET.
-  • Extending the Benchmarking by further showing the capabilities of BOUQuET, e.g.
-increasing the evaluation of linguistic signals over its alternatives (FLORES-200, NTREX-128, and NLLB-MD).
-For this, we propose to explore the performance in different domains and see if we get difference in the performance ranking of translation systems.
-Acknowledgments
-We would like to extend our sincere gratitude to Luis Cadavid, Radhi Datla, Suci Fitriany, Christine Haugen, Stephanie Mauro, Carissa Ikka Pardamean, Gauri Shanware, and Jyoti Singh for their invaluable contributions to the work presented in this paper.
+We have shown consistent gains in domain diversity in two different metrics while keeping complexity lower than its competitors.
+The latter is particularly relevant to simplify the translation for non-experts that may join the open initiative.
+We also provide MT results for the 8 languages in which Source-BOUQuET has been created.
+Although BOUQuET is currently totally completed for 55 languages (see list 6), this number is only a fraction of the language coverage ambition that we are pursuing by launching the open initiative for community efforts.
+Please join us in making Universal Quality Evaluation in Translation available in any language.
+Beyond increasing in number of languages, BOUQuET is actively evolving, and we are currently working on designing quality control for each of the contributions and adding new languages to the incremental releases of BOUQuET and extending the benchmarking by further showing the capabilities of BOUQuET, e.g.
+increasing the evaluation of linguistic signals over its alternatives.
+Limitations and Ethical Considerations
+The BOUQuET dataset is still limited in the number of languages and translations.The benchmarking is quite complete (4 datasets comparison, 14 models and 2 metrics) but it can also be extended in several axes (linguistic analysis).
+However, the entire purpose of this work is to describe the dataset and open-initiative, while providing a minimal benchmarking.
+Authors expect the community to extend the benchmarking by further using this dataset for further exploration.
+Creators and commissioned translation’s annotators are paid a fair rate.
 Appendix A Specific guidance for paragraph and sentence creation
 A.1 Overview
-The Bouquet-source dataset needs to include 250 sentences in each of its source languages.
-This means that each linguist needs to create (i.e., create from scratch, not copy; see Section 2.4 above) 250 original sentences.
-These sentences must be:
+The Bouquet-source dataset comprises 250 unique sentences in each of its source languages.
+This means that each linguist created (i.e., wrote from scratch, did not copy; see Section 2.4 above) 250 original sentences.
+These sentences were requested to be:
   • Organized in logically structured paragraphs (see the Paragraphs section below)
@@
 A.2 Paragraphs
-The linguist will receive a template in the form of a spreadsheet, in which paragraph structures are designed and laid out.
-The template specifies the exact number of paragraphs and the exact number of sentences for each of the paragraphs.
-Each paragraph is given a unique paragraph ID (e.g., P01, P02, P15).
-Each sentence within each paragraph is also given a serial, non-unique ID (e.g., S1, S2, S3).
-The existing structure must not be modified in any way.
+The linguist received a template in the form of a spreadsheet, in which paragraph structures were designed and laid out.
+The template specified the exact number of paragraphs and the exact number of sentences for each of the paragraphs.
+Each paragraph was given a unique paragraph ID (e.g., P01, P02, P15).
+Each sentence within each paragraph was also given a serial, non-unique ID (e.g., S1, S2, S3).
 A.3 Domains
-The template is divided into 8 domains:
+The template was divided into 8 domains:
   1. 1.
@@
     Miscellaneous (address to a nation, disaster response, etc.)
-You must produce the set number of sentences for each of the domains; the domain / paragraph / sentence structure of the template cannot be changed.
+The creators had to produce the set number of sentences for each of the domains; the structure of the template (domain / paragraph / sentence) could not be changed.
 A.4 Language Register Information
-When creating sentences, make sure that the register of language being used is representative of the most expected and appropriate register for the situation.
-We understand that several registers could be possible; please use discretion when selecting a register, while making sure that the chosen register is among the most expected and appropriate.
-To help you make a determination, we define 3 main functional areas of language register:
+When creating sentences, the creators had to make sure that the register of language being used was representative of the most expected and appropriate register for the situation.
+When several registers were possible, the creators were asked to use discretion when selecting a register, while making sure that the chosen register was among the most expected and appropriate.
+To help them make a determination, we defined 3 main functional areas of language register:
   • Connectedness: What type of connection do language users who initiate the text have with other language users?
@@
 A.6 Violating Content
-While creating sentences, please be sure to avoid inserting violating content. Violating content is language that can fall under one (or more) of the below categories:
+While creating sentences, the creators were asked to avoid inserting violating content. Violating content is language that can fall under one (or more) of the below categories:
   • Toxicity
@@
 A.7 Step-by-Step Description of Tasks
-Please use the provided template to follow this step-by-step description.
+Please refer to Table 5 for the step-by-step description of the tasks.
 Column A: Lang-ID This column should have the same 3-lowercase-letter code representing the source language of the sentences
@@
 Column T: P-Nchars This represents a count of the number of characters in the current paragraph.
-Column U: Creator_Translator-ID Please insert your ID here, if it isn’t prepopulated.
+Column U: Creator_Translator-ID Please insert your ID here, if it isn’t pre-populated.
+Table 5: Step-by-step guidance.
 A.8 Additional Guidance on Domain-Specific Content
 Dialogues, especially those inserted in long creative writing (such as novels), often include the name of the speaker or a cue mark (e.g., — ), and sometimes quotation marks.
-When creating sentences for conversations, please feel free to invent names for speakers or to label speaker turns (e.g., A, B).
-Please place the names or speaker reference in markup tags, similarly to this: <Name:>or <A:>.
-Emojis: There are emojis frequently in some social media and messaging domains.
-There should therefore be a representation, albeit very limited but the point isn’t to know how the system translates emojis because there are no real agreed ways to translate them across hundreds of languages.
-Social media comments: Okay if the structure is flattened. Okay to have some tags but not absolutely necessary and no need to insert them everywhere.
-Disfluencies in informal conversations?: If they are representative of conversations and they can be translated (i.e., there is some consensus on how to write them in the language — ah, oh, um).
-Appendix B Contribution Statement
-Pierre Andrews: Open initiative; Supervision/Leadership; core; tool deployment and design
-Mikel Artetxe: Benchmarking; Supervision; support
-Mariano Coria Meglioli: Dataset; core; statistical analysis
-Marta R. Costa-jussa: Dataset, Benchmarking, Tooling Supervision/Leadership; core; corresponding author; research direction, paper editor
-Joe Chuang: Dataset; core; source-bouquet creator
-David Dale: Benchmarking; Supervision/Leadership; core; evaluation and experimental design
-Cynthia Gao: Dataset; support
-Jean Maillard: Open initiative; core, licensing
-Alex Mourachko: Management; support
-Christophe Ropers: Dataset and Open initiative; Supervision/Leadership; core; dataset design and creation, guidelines design and creation, paper editor
-Safiyyah Saleem: Management; support
-Eduardo Sánchez: Benchmarking; core; experiments deployment
-Yiannis Tsiamas: Benchmarking; support
-Arina Turkatenko: Dataset; core; dataset creation
-Albert Ventayol-Boada: Dataset; core; dataset creation
-Shireen Yates: Management; support
+When creating sentences for conversations, the creators were let free to invent names for speakers or to label speaker turns (e.g., A, B); they were also asked to place the names or speaker reference in markup tags, similarly to this: <Name:>or <A:>.
+Emojis: As there are emojis frequently in some social media and messaging domains, some representation was also expected from the creators.
+However, the creators were asked to keep this representation very limited, as there are no real agreed ways to translate them across hundreds of languages.
+Social media comments: The creators were told that they could keep the structure of those comments flat, and that including tags was not absolutely necessary, though it was permitted (even expected).
+Disfluencies in informal conversations: Disfluencies were permitted provided they were representative of conversations and they could be translated (i.e., there is some consensus on how to write them in the language — ah, oh, um).
+Appendix B Registers Details
+Figure 5: Register functional areas and breakdowns within each functional areas and their representations across domains.
+We provide non-exclusive options for each of the 3 functional areas that characterize registers described in Section 3.2 and mentioned in Figure 5.
+By non-exclusive, we mean that a domain may be characterized by more than one option.
+The functional area / option breakdown can be described as follows (the bold lowercase letters in square brackets represent a unique identifier for each option):
+Connectedness
+  • Impersonal [i]: For example a text written for the purpose of giving definitions or explanations with no specific readership in mind; typically written in the third person only (e.g., a contract).
+  • Non-directional [n]: A text written with a readership in mind but that doesn’t address the readership specifically (e.g., an author recounting a story)
+  • Uni-directional [u]: A text addressing a readership who either cannot respond or is asked to refrain from responding at a given time (e.g., the transcription of a presentation, such as a TED Talk)
+  • Multi-directional [m]: A text addressing a readership who can respond (e.g., SMS, DM) or representing the transcription of a dialogue involving 2 or more language users.
+Preparedness
+  • Reactive (spontaneous) [r]: The production is immediate either because it needs to be or because the user wants it to be
+  • Improvised (coached) [i]: The production appears spontaneous but takes place after a period of general training or coaching (e.g., spokespeople who answer questions live but have had time to prepare and choose vocabulary to use or to avoid)
+  • Rehearsed (extemporaneous) [e]: The production is live but its overall structure has been carefully crafted and rehearsed (e.g., transcriptions of 20-minute presentations or speeches that aren’t fully scripted and given from notes).
+  • Scripted (declaimed) [s]: The production may or may not be live and has been fully scripted (e.g., transcriptions of speeches used in teleprompters)
+  • Composed (frozen) [c]: The production is completely offline, and goes through iterations of reviewing and editing (e.g., the text of a novel).
+Social differential
+  • Equal (known) [k]: The readership or addressees are known to be peers; this can include a very informal or colloquial attitude
+  • Equal (assumed) [a]: The readership or addressees are not known but assumed to be peers; this can include a casual or informal attitude but likely excludes a very colloquial one
+  • Higher-to-lower [h]: The readership or addressees are considered to be at a lower social level than the producer (e.g., the producer is arrogant or assumes a position of higher authority)
+  • Lower-to-higher [l]: The readership or addressees are considered to be at a higher social level than the producer (e.g., the producer wants to express deference, respect, or admiration)
+Appendix C Quality Checks Details in Multi-way extension
+In order to make sure that the BOUQuET contextual information was taken into accound while translating BOUQuET, we used the following QA strategies:
+  1. 1.
+    Checking the correct co-referencing.
+The Bouquet dataset is a representation of natural language, and the usage of personal and possessive pronouns as a substitute for the nouns is a typical occurrence.
+If the internal co-referencing in the paragraph is broken (the wrong pronoun is used or the noun is repeated where the noun should be), it indicates that the paragraph was treated as a collection of sentences not linked to each other, rather than a paragraph of text.
+  2. 2.
+    Checking the lexical consistency.
+We made sure to check that vocabulary used to translate word denoting objects or events is appropriate in tone, style and register and is used consistently throughout each paragraph.
+For example, when checking, we found out that translations from Indonesian into Russian did not keep consistency for “potato fritters” (“perkedel kentang”), using three different ways to translate it in P-292.
+We later applied the necessary corrections.
+  3. 3.
+    Checking the grammatical consistency.
+Since the Bouquet dataset contains examples of different domains, we needed to check whether the verb tenses and syntax were appropriate for a given domain and used consistently throughout each paragraph.
+For example, when checking translated into German paragraphs which imitate fiction narration, we made sure that German Pratäritum tense is used appropriately, not Perfekt.
+  4. 4.
+    Checking the special symbols such as emojis and numbers.
+Appendix D Priority Languages
+Table 6 shows the languages in which BOUQuET exists at the time of submission of this paper (May 2025).
+ISO 639-3 ISO 15924 Language Family Subgroup Class
+arz Arab Egyptian Arabic Afro-Asiatic Central Semitic Pivot
+(+ arb +Modern Stan. Arabic)
+arz Latn Romanized Egyptian Arabic Afro-Asiatic Semitic P1-HR
+aar Latn Afar Afro-Asiatic Cushitic P1-LR
+agr Latn Aguaruna Chicham P1-LR
+ami Latn Amis Austronesian East Formosan P1-LR
+ben Beng Bengali Indo-European Indo-Aryan P1-HR
+cmn Hans Mandarin Chinese Sino-Tibetan Sinitic Pivot
+ces Latn Czech Indo-European Balto-Slavic P1-HR
+crk Cans Plains Cree Algic Algonquian P1-LR
+deu Latn German Indo-European West Germanic Pivot
+dje Arab, Latn Zarma Songhay Eastern Songhay P1-LR
+ell Grek Modern Greek Indo-European Hellenic P1-HR
+fra Latn French Indo-European Italic Pivot
+gaz Latn West Central Oromo Afro-Asiatic Cushitic P1-LR
+gil Latn Gilbertese Austronesian Micronesian P1-LR
+guc Latn Wayuu Arawakan Caribbean Arawakan P1-LR
+hin Deva Hindi Indo-European Indo-Aryan Pivot
+hin Latn Romanized Hindi Indo-European Indo-Aryan P1-HR
+hrv Latn Croatian Indo-European Balto-Slavic P1-HR
+hun Latn Hungarian Uralic Hungaric P1-HR
+ind Latn Indonesian Austronesian Malayic Pivot
+ita Latn Italian Indo-European Italic P1-HR
+jav Latn Javanese Austronesian Javanesic P1-HR
+jpn Jpan Japanese Japonic P1-HR
+kaa Cyrl Karakalpak Turkic Kipchak P1-LR
+kal Latn Kalaallisut Eskimo-Aleut Eskimo P1-LR
+khm Khmr Central Khmer Austroasiatic Mon-Khmer P1-HR
+kor Kore Korean Korean Koreanic P1-HR
+kru Deva Kurukh Dravidian North Dravidian P1-LR
+lij Latn Ligurian Indo-European Italic P1-LR
+lin Latn Kinshasa Lingala Atlantic-Congo Central West. Bantu P1-LR
+mya Mymr Burmese Sino-Tibetan Burmo-Qiangic P1-LR
+nld Latn Standard Dutch Indo-European West Germanic P1-HR
+pes Arab Western Persian Indo-European Iranian P1-HR
+pol Latn Polish Indo-European Balto-Slavic P1-HR
+rus Cyrl Russian Indo-European Balto-Slavic Pivot
+ron Latn Romanian Indo-European Italic P1-HR
+sba Latn Ngambay Central Sudanic Sara-Bongo-Bagirmi P1-LR
+spa Latn Spanish Indo-European Italic Pivot
+por Latn Portuguese (Brazilian) Indo-European Italic P1-HR
+swe Latn Swedish Indo-European North Germanic P1-HR
+swh Latn Coastal Swahili Atlantic-Congo N.E. Coastal Bantu P1-HR
+tha Thai Thai Tai-Kadai Southwestern Tai P1-HR
+tir Ethi Tigrinya Afro-Asiatic Semitic P1-LR
+tgl Latn Tagalog Austronesian Greater Central Philippine P1-HR
+tur Latn Turkish Turkic Oghuz P1-HR
+ukr Cyrl Ukrainian Indo-European Balto-Slavic P1-HR
+urd Arab Urdu Indo-European Indo-Aryan P1-HR
+vie Latn Vietnamese Austroasiatic Vietic P1-HR
+yor Latn Yoruba Atlantic-Congo Defoid P1-LR
+zlm Latn Colloquial Malay Austronesian Malayic P1-HR
++zsm + Standard Malay
+Table 6: Source-BOUQuET Languages (Pivot) and Priority languages (P) both high-resource (HR) and low-resource (LR) included in BOUQuET at the time of submission.
+Note that these languages have been commissioned, we do not include updates in annotations collected from the open-initiative, which we will include in later versions of the paper.
+Appendix E Dataset Examples
+Table 7 reports the sentences with highest similarity score computed with cosine distance of SONAR vectors across all 2,000 Source-BOUQuET English sentences.
+cosinedist lang-A lang-B Domain-A Text-A UNIQID-A DomainB Text-B UNIQID-B
+0.19 ind arz conversation What time do we meet? P304-S4 conversation when will we meet? P017-S1
+0.20 fra cmn web misc. About us P220-S1 web misc. About our team P098-S1
+0.22 rus deu conversation <<<B:>>> Which one? P363-S2 conversation <<<B:>>> When and where? P134-S2
+0.24 rus fra conversation <<<B:>>> Nah, I am sick P360-S2 conversation <<<B:>>>You’re sick? P185-S4
+0.25 rus fra conversation <<<A:>>> You know what I mean! P362-S4 conversation <<<A:>>>Did you hear? P183-S1
+0.28 fra arz web misc. Send us your résumé and motivation letter at the below address. P215-S6 web misc. Please sendyour CV with letters of recommendation to this email address P043-S5
+0.28 spa rus comments <<<B:>>> WHAT IS THIS??? P443-S2 conversation <<<B:>>> What do you mean? P362-S2
+0.29 rus fra conversation <<<A:>>> Get well soon P360-S3 conversation <<<A:>>>Not doing very well. P185-S3
+0.29 rus fra conversation <<<B:>>> What do you mean? P362-S2 conversation <<<A:>>>Did you hear? P183-S1
+0.29 rus fra conversation <<<B:>>> What do you mean? P362-S2 conversation <<<B:>>>You’re sick? P185-S4
+0.29 rus fra conversation <<<B:>>> Nothing is working for me. P366-S2 conversation <<<A:>>>Not doing very well. P185-S3
+Table 7: Source-BOUQuET sentences with closest similarity score (cosine distance lower than 0.3)
+Table 8 shows complete entries examples of the Source-BOUQuET dataset.
+LangID Domain Subdomain PID SID Sentence English Linguistic label Reg.
+spa_Latn conversation text message chain P417 S1 <<<Guillermo:>:absent:>: > Habéis cenado ya? <<<Guillermo:>:absent:>: > Have you had dinner already? word:named-entity mrk
+spa_Latn conversation text message chain P417 S2 <<<Jaime:>:absent:>: > No, estábamos pensando en salir ahora, te apuntas?
+<<<Jaime:>:absent:>: > No, we were thinking about going out now.
+Are you in?
+word:named-entity mrk
+spa_Latn conversation text message chain P417 S3 <<<Guillermo:>:absent:>: > Sí, me estoy muriendo de hambre.
+<<<Guillermo:>:absent:>: > Yes, I’m starving.
+word:named-entity, miscellaneous:collocation mrk
+spa_Latn conversation text message chain P417 S4 <<<Jaime:>:absent:>: > Guai, salimos en cinco, te esperamos en la parada del metro.
+<<<Jaime:>:absent:>: > Cool, we’re leaving in five, we’ll wait for you at the metro station.
+word:named-entity, word:slang mrk
+spa_Latn conversation text message chain P417 S5 <<<Guillermo:>:absent:>: > Perfecto, me cambio y salgo. <<<Guillermo:>:absent:>: > Perfect, I’ll change and head out. word:named-entity mrk
+fra_Latn social posts Integrity P204 S1 Choses que j’aurais aimé savoir plus tôt Things I wish I had known earlier sentence:fragment usa
+fra_Latn social posts Integrity P204 S2 Si tu ne prends pas de décision pour toi-même, d’autres les prendront pour toi.
+If you don’t make decisions for yourself, others will take them for you.
+word:impersonal-pronoun usa
+fra_Latn social posts Integrity P204 S3 Quand on te submerge de généralités, demande plusieurs exemples spécifiques.
+When you are getting submerged in generalities, request several specific examples.
+word:impersonal-pronoun usa
+ind_Latn narration Folklore / Fable P310 S1 Pada suatu masa, hiduplah sepasang suami istri di sebuah pedesaan.
+Once upon a time, there lived a husband and wife in a village.
+Third person, impersonal, narration ica
+ind_Latn narration Folklore / Fable P310 S2 Mereka belum juga dikarunia anak setelah sekian lama menikah.
+They have not yet been blessed with children after being married for so long.
+Third person, impersonal, narration ica
+ind_Latn narration Folklore / Fable P310 S3 Keduanya bermimpi bahwa mereka harus menanam timun, jika mereka ingin memiliki anak.
+Both of them dreamed that they had to plan cucumbers, if they wanted to have a child.
+Third person, impersonal, narration ica
+ind_Latn narration Folklore / Fable P310 S4 Kemudian ditanamlah timun-timun itu. Then they planted the cucumbers. Third person, impersonal, narration ica
+Table 8: BOUQuET examples including main fields
+Appendix F Domain representation details
+Figures 6 shows the domain representation and overlap across datasets.
+Figure 6: Domain representation and overlap across FLORES+ (left), NTREX-128 (middle), NLLB-MD (right)(in blue) with diverse domains datasets (in grey) and BOUQuET (in red).
+Appendix G Detailed results
+Src-lang arz-Arab cmn-Hans deu-Latn eng-Latn fra-Latn hin-Deva ind-Latn rus-Cyrl spa-Latn
+com metx com metx com metx com metx com metx com metx com metx com metx com metx
+nllb-3B 0.59 2.31 0.66 1.39 0.72 2.39 0.75 1.86 0.69 2.11 0.61 2.33 0.69 2.03 0.67 2.33 0.72 2.13
+aya101-13B 0.59 2.23 0.65 1.32 0.71 2.31 0.75 1.76 0.67 2.07 0.6 2.32 0.69 1.88 0.67 2.26 0.71 1.99
+aya-e-8B 0.6 2.04 0.68 1.09 0.73 2.06 0.77 1.47 0.7 1.77 0.62 1.94 0.7 1.69 0.69 1.94 0.73 1.76
+babel-9B 0.57 2.78 0.66 1.61 0.72 2.58 0.75 1.92 0.68 2.47 0.59 2.54 0.69 2.16 0.67 2.47 0.71 2.46
+cohere-r7B 0.58 2.38 0.66 1.28 0.72 2.45 0.75 1.87 0.66 2.52 0.61 2.18 0.7 1.86 0.66 2.65 0.71 2.16
+eurollm-9B 0.58 2.55 0.66 1.59 0.71 2.68 0.75 2.04 0.68 2.38 0.61 2.41 0.66 2.37 0.67 2.58 0.71 2.36
+madlad-10B 0.52 3.71 0.62 1.8 0.66 3.23 0.73 2.0 0.64 2.61 0.58 2.75 0.62 3.06 0.63 3.08 0.69 2.39
+madlad-3B 0.51 3.93 0.62 1.76 0.63 3.51 0.72 2.26 0.63 2.82 0.59 2.63 0.61 3.25 0.63 3.1 0.7 2.4
+mistral-7B 0.44 5.32 0.54 3.53 0.59 4.52 0.6 3.83 0.55 4.2 0.46 4.95 0.59 3.81 0.56 4.31 0.59 4.11
+qwen-7B 0.5 3.83 0.58 2.55 0.63 3.49 0.68 2.79 0.6 3.25 0.52 3.45 0.61 3.16 0.59 3.47 0.62 3.29
+Llama-3.1-8B 0.56 2.77 0.64 1.58 0.7 2.63 0.75 1.88 0.66 2.46 0.58 2.77 0.68 2.24 0.66 2.52 0.7 2.38
+Llama3.2-3B 0.46 4.88 0.58 2.56 0.63 3.86 0.67 3.02 0.58 3.77 0.53 3.78 0.61 3.35 0.59 3.46 0.63 3.62
+Llama3.3-70B 0.61 1.96 0.68 1.18 0.74 2.14 0.77 1.62 0.71 1.94 0.62 2.1 0.71 1.73 0.69 2.05 0.74 1.91
+Tower-7B 0.4 6.07 0.62 1.5 0.63 3.23 0.66 3.48 0.56 4.21 0.49 4.42 0.62 3.26 0.6 3.61 0.64 3.46
+Trg-lang arz-Arab cmn-Hans deu-Latn eng-Latn fra-Latn hin-Deva ind-Latn rus-Cyrl spa-Latn
+com metx com metx com metx com metx com metx com metx com metx com metx com metx
+NLLB-3B 0.62 3.08 0.59 2.99 0.71 0.99 0.76 2.22 0.69 2.01 0.61 2.55 0.7 1.53 0.71 1.77 0.72 1.76
+aya101-13B 0.64 2.61 0.63 1.82 0.69 1.07 0.76 2.26 0.68 2.14 0.56 2.94 0.69 1.55 0.69 1.85 0.7 1.89
+aya-e-8B 0.7 1.9 0.65 1.81 0.72 0.85 0.77 2.01 0.71 1.82 0.53 3.02 0.71 1.3 0.72 1.47 0.73 1.59
+babel-9B 0.64 3.06 0.66 1.96 0.68 1.31 0.76 2.1 0.69 2.06 0.53 3.74 0.68 2.42 0.69 2.42 0.71 1.91
+cohere-7B 0.68 2.38 0.65 1.87 0.7 1.0 0.76 2.13 0.69 1.95 0.51 3.99 0.66 2.08 0.68 2.24 0.72 1.72
+eurollm-9B 0.7 1.99 0.66 1.65 0.72 0.89 0.77 2.13 0.7 1.87 0.6 2.65 0.45 6.53 0.72 1.57 0.72 1.67
+madlad-10B 0.65 2.67 0.59 2.61 0.67 1.64 0.75 2.47 0.66 2.71 0.41 5.12 0.62 2.64 0.66 2.66 0.7 2.1
+madlad-3B 0.65 2.8 0.58 2.77 0.66 1.73 0.73 2.79 0.64 2.81 0.42 5.1 0.62 2.64 0.65 2.83 0.69 2.2
+mistral-7B 0.32 8.73 0.55 3.12 0.62 1.85 0.73 2.7 0.62 2.96 0.3 8.42 0.52 4.52 0.61 3.38 0.65 2.89
+qwen-7B 0.53 4.74 0.64 2.09 0.63 1.76 0.72 2.39 0.64 2.63 0.25 7.38 0.64 2.65 0.61 3.12 0.66 2.49
+Llama3.1-8B 0.54 4.68 0.64 1.91 0.69 1.17 0.76 2.25 0.68 2.17 0.57 3.05 0.68 1.81 0.68 2.18 0.7 2.01
+Llama3.2-3B 0.43 6.84 0.55 3.11 0.63 1.76 0.73 2.62 0.62 2.86 0.49 4.5 0.62 2.7 0.54 5.27 0.66 2.66
+Llama3.3-70B 0.6 3.32 0.68 1.62 0.73 0.85 0.77 2.04 0.71 1.84 0.63 2.41 0.72 1.35 0.72 1.55 0.73 1.64
+Tower-7B 0.3 7.96 0.61 2.33 0.67 1.44 0.74 2.74 0.66 2.6 0.41 6.54 0.51 4.62 0.66 2.52 0.68 2.5
+Table 9: Averaged results on CometKiwi (COM) and MetricX (etx) at the sentence-level from 9 BOUQuET languages (top) and into (bottom).
+Best results are in bold (before rounding to 2 decimals).
+Best results on CometKiwi tend to be with Llama-3.3-70B (the largest model) and best results in MetricX tend to be with Aya-expanse-8B.
+Best direction is from and into English .